OpenStudy (anonymous):

Antonio is seven less than four times his daughter's age. Antonio is 37; how old is his daughter?

OpenStudy (mertsj):

d = daughter's age 4d-7= antonio's age 4d-7=37

OpenStudy (anonymous):

hard to understand..

OpenStudy (anonymous):

okay guys the choices are 11 12 18 6

OpenStudy (anonymous):

11

OpenStudy (anonymous):

7-4d=37 -4d=37-7 -4d=20 -4d/4=20/-4d d=-5

OpenStudy (anonymous):

good @Luis Rivera . you make things easier.